OpenStudy (anonymous):

using the discriminant, determine the number and type of solutions for this equation 5x^2-10x-2=0

OpenStudy (anonymous):

\[5x^{2}-10x-2\] Discriminant= \[b ^{2}-4ac\] =\[(-10)^{2}- 4 (-2)(5)\] \[=100+40=140\] since the discriminant is positive \[ \sqrt{D}=2\sqrt{35}\] and it's not a perfect square thus it will have 2 distinct Real roots. hope u get it :-)